# SC-Bench: A Large-Scale Dataset for Smart Contract Auditing

### v1

@@ Abstract
There is a huge demand to ensure the compliance of smart contracts listed on blockchain platforms to safety and economic standards. Today, manual efforts in the form of auditing are commonly used to achieve this goal. ML-based automated techniques have the promise to alleviate human efforts and the resulting monetary costs. However, unlike other domains where ML techniques have had huge successes, no systematic ML techniques have been proposed or applied to smart contract auditing. We present SC-Bench, the first dataset for automated smart-contract auditing research. SC-Bench consists of 5,377 real-world smart contracts running on Ethereum, a widely used blockchain platform, and 15,975 violations of standards on Ehereum called ERCs. Out of these violations, 139 are real violations programmers made. The remaining are errors we systematically injected to reflect the violations of different ERC rules. We evaluate SC-Bench using GPT-4 by prompting it with both the contracts and ERC rules. In addition, we manually identify each violated rule and the corresponding code site (i.e., oracle) and prompt GPT-4 with the information asking for a True-or-False question. Our results show that without the oracle, GPT-4 can only detect 0.9% violations, and with the oracle, it detects 22.9% violations. These results show the potential room for improvement in ML-based techniques for smart-contract auditing.
1 Introduction
@@
Ethereum [35, 13] is a decentralized, open-source blockchain platform that has become the de facto for running decentralized applications like smart contracts [14, 17]. To govern smart contracts running on Ethereum, a set of formal standards called Request for Comments, or ERCs, have been developed [31]. For example, the ERC20 standard [33] defines a common set of rules for fungible tokens — digital assets that are interchangeable with one another. ERCs play a crucial role in the Ethereum ecosystem by providing a common set of rules and specifications that developers can follow when implementing smart contracts. The violation of an ERC could result in interoperability issues, security vulnerabilities, and financial loss. Additionally, ERC violations could result in the de-listing of tokens from exchanges, as many exchanges have listing requirements of following ERC standards [15].
Despite the importance of following ERC rules, it is hard for developers to always do so, as they need a comprehensive understanding of all ERC requirements and their contract code. ERC standards include a huge number of rules, and the number keeps growing as standards are developed over time. For just the three ERC standards involved in this work, there are 153 rules. ERC rules are described in different ways, with some being code declarations and others being natural languages. Meanwhile, contract implementations are often complex as well. One contract and its dependent code usually contain hundreds to thousands of lines of source code in multiple files. Some code details may be obscured within intricate caller-callee relationships, while others may involve objects and functionalities written by different programmers. Different rules apply to different parts of a smart contract, and even a small code site could involve multiple rules. These complexities in ERC rules and smart contracts make it extremely hard for programmers to check for ERC violations manually. As a result, ERC rule violations widely exist in real-world smart contracts [8].
@@
To detect ERC violations, today’s common practice heavily relies on human efforts. Automated, program-analysis-based checkers for ERC rules do exist [9, 24], but they fail to detect complex violations because of many ERC rules’ non-structured, natural-language-based definitions. As a result, smart contracts commonly undergo human auditing, provided by specialized services with security experts [6, 28, 25, 3, 20, 1, 8]. Auditing services are not only costly but also slow. For example, we examined the history of 30 smart contracts that were submitted for manual auditing on a platform called Ethereum Commonwealth Security Department. We found that each contract only has an average of 260 lines of code but is audited for ten days with an estimated cost of \$500. Clearly, manual auditing is not a scalable approach for smart contract development.
@@
A promising, scalable approach for automated smart-contract auditing is to leverage large language models (LLMs), both because of LLMs’ success in domains like program generation [30] and bug fixing [21, 38] and because of a fair amount of ERC rules’ natural-language descriptions. To develop LLM-based techniques (e.g., fine tuning, in-context learning, few-shot learning, etc.) for smart contract auditing, an important step is to construct quality datasets. Unfortunately, no smart contract datasets exist that can be used for ERC rule checking and fixing. Traditional program bug datasets [29, 19] cannot be used for smart-contract auditing, as unlike ERC rules, compiler errors and program runtime errors have well-defined, structured definitions.
@@
To drive research and practices in smart-contract auditing and to assist real users with their auditing tasks, we release SC-Bench, the first dataset of real-world smart contracts and their ERC-rule violations. SC-Bench consists of 15975 ERC violations and 5377 real-world smart contracts, collected from etherscan.io [18], polygonscan.com [26], and Ethereum Commonwealth Security Department [8]. 145 violations from 30 contracts are real-world ERC violations we collect and inspect. As real violations are rarely published, we build program analysis techniques and inject 15836 violations into 5347 contracts according to 88 ERC rules.
@@
We evaluate SC-Bench using GPT-4 with two methodologies. The first prompts GPT-4 with the contract to be inspected and the full ERC rule sets. Our results show that GPT-4 only detects 29% of real violations and 0.6% of injected violations. To test how GPT-4 could potentially detect violations given oracle information, we manually identify the violated ERC rule(s) and violating code sites for each contract. We then prompt GPT-4 with this precise information and ask it a True-or-False question. GPT-4 successfully detects 42.8% of real violations and 22.8% injected violations. The increase in detecting indicates the potential room for ML-based techniques to improve smart contract automated auditing.
@@
Overall, SC-Bench goes beyond being a key contribution to Ethereum smart contract auditing. Other software auditing and verification tasks, such as security defenses, can potentially use SC-Bench’s contract dataset, its violation dataset, or its methodologies for evaluation. We have released SC-Bench and all our source code at https://github.com/system-pclub/SC-Bench.
2 Background
@@
2.1 Ethereum and Smart Contracts
@@
2.2 Ethereum Request for Comment (ERC)
@@ 2.3 ERC Rule Violations
2.4 Today’s Auditing Practices
The common practice for detecting ERC rule violations today relies on manual auditing, often provided by paid services [6, 28, 25, 3, 20, 1, 8]. One such service is the Ethereum Commonwealth Security Department [8], where users submit smart contracts for auditing by filing a GitHub issue. The service then manually audits the submitted contracts and provides feedback through the issue. To reduce the manual workload and associated costs, some automated tools have been developed using static program analysis. For example, Slither offers specific checkers (i.e., slither-check-erc [9]) that verify whether a given contract complies with the corresponding ERC standards for 11 ERCs. However, these tools have limited functionality. They primarily focus on ensuring the presence of required functions and events, confirming that these elements are correctly declared, and verifying that functions trigger the necessary events. Unfortunately, they are unable to check more advanced conditions, such as verifying whether the message caller has enough privilege to transfer tokens. To our knowledge, no machine learning-based automated auditing techniques have been proposed yet. We hope that the release of SC-Bench can foster a new line of research in this regard.
Table 1: How ERC rules are distributed across different error-injection methods. (“Uncovered”: rules whose violations cannot be injected by our designed error-injection methods.)
@@ 2.4 Today’s Auditing Practices
3 SC-Bench
3.1 Construction
We collect real-world smart contracts from Ethereum Commonwealth Security Department [8], etherscan.io [18], and polygonscan.com [26]. As the purpose of SC-Bench is for evaluating automated ERC-rule checking, we include ERC rule violations of these collected contracts in SC-Bench using two approaches. First, we manually inspect a set of real smart contracts and identify all their ERC violations. This process is time-consuming, resulting in a limited number of violations. To address this, we perform automated error injection into a large number of real smart contracts, significantly increasing the number of violations. These two types of violations serve to validate each other and help ensure that the evaluated techniques demonstrate consistent performance.
Manual Inspection. We manually analyzed 30 ERC20 contracts obtained from the Ethereum Commonwealth Security Department [8], selecting the most recent 30 audit requests that meet the following criteria: 1) they contain Solidity source code, 2) they have been approved by Solidity developers (as indicated by the “approved” tag), 3) they exhibit ERC rule violations, and 4) all contract and related Solidity code is within the same contract file. On average, each contract file contains 260.9 lines of Solidity code. Through detailed examination, we identify 139 ERC rule violations. Of these, 27 violations present a clear method for exploitation that can lead to financial losses, and we classify them as having a high-security impact. Another 48 violations result from incorrect implementation of required functionalities, but there is no apparent way to exploit these for financial gain. Therefore, we consider them to have a medium-security impact. The remaining 64 violations involve failures in generating necessary logs and are categorized as having a low-security impact.
Error Injection. To augment the limited manual-inspected real violations, we perform error injection to 5,347 real-world smart contract source code using the following methodology. We design six error injection methods corresponding to 85 ERC rules, resulting in a total of 15,836 ERC violations. Table 1 shows the number of rules each method covers.
We focus on contracts in three ERC standards, ERC20 [33], ERC721 [12], and ERC1155 [27], for two reasons. First, these ERCs are significant and have numerous crucial financial applications. For instance, there are over 450,000 ERC20 tokens on the Ethereum platform [5], many of which have market capitalizations exceeding \$1 billion (e.g., USDT [37], SHIB [36], Binance USD [34]). Second, these three standards are among the most mature ERCs and inspire subsequent ERCs, making their rules representative of rules in other ERCs. As shown in Table 1, these three ERCs specify 132 rules. Our error injection methods can cover 85 of them. We do not include the rest either because they are not clearly specified (e.g., “throw if any other error” in ERC1155) or because they require more complex static analysis or injection methods.
Table 2: How many errors are injected by different error-injection methods.
# of Violations # of Contracts
@@ 3.1 Construction
We collect contracts in these three ERCs from etherscan.io [18] and polygonscan.com [26]. These two platforms are the most popular analytics platforms for Ethereum and its sidechain Polygon [22], respectively. As shown in Table 2, we collected 5,211 contracts that are supposed to target ERC20, 110 for ERC721, and 26 for ERC1155111We mainly rely on a contract’s name or the name of a base contract inherited by the contract to determine whether it implements a particular ERC. . On average, each of these contracts contains 477.5 lines of code. We collect more contracts targeting ERC20 than ERC721 and ERC1155 contracts due to the significantly higher prevalence of ERC20 in practice.
@@
We randomly inject one to three errors in each contract based on the fit of rules to the contract to construct ERC violations. To inject an error, we first convert each input smart contract source-code file into its abstract syntax tree (AST). We then randomly select a rule and apply the corresponding error injection to the AST. After modifying the AST, we convert it back into source code. We perform error injections on ASTs because it is easier to conduct static program analysis and modify code on ASTs than to work directly with the source code. Table 2 shows the number of errors injected by each method. Finally, we verify that the modified source code compiles without errors using the Solidity compiler. Below, we discuss each error injection method in detail.
@@
Violations of Condition-Check Rules. ERCs mandate certain public functions to perform condition checks on their input parameters or callers’ addresses (i.e., msg.sender). Sometimes, these checks validate the input parameters. For example, ERC721 disallows the input parameter of function ownerOf() to be zero. More crucially, these checks ensure the caller has the permission to perform an action. For instance, ERC20’s transfer() function sends tokens in the amount specified by its second parameter from the caller’s account to the receiver, whose address is the first parameter. ERC20 requires verifying that the caller has enough tokens (an amount greater than or equal to the second parameter) and throwing an exception if this condition is not met. Since many of these rules are implemented with require statements, to inject an error of this type into a function, we remove all require statements that include the parameter required to be checked by the rule as part of their conditions.
@@
Violations of Return-Value Rules. ERCs explicitly define how return values should be computed for certain function call APIs. For example, ERC20 requires that allowance() returns the token amount that the first parameter allows the second parameter to withdraw. As return-value rules apply to only four data types, integer, Boolean, address, and string, we design four distinct error-injection methods for each of them. We add a random integer to an integer return value, flip a Boolean return value, replace an address return with a random address, and change a string return value to an empty string if it is not already empty or replace an empty return string with a random string.
@@
Violations of Value-Update Rules. Fields or state variables in a contract represent its state. Some ERC rules specify how a public function should update a state variable. For instance, contracts following ERC20 rules use a two-dimensional mapping to record the number of tokens the first key allows the second key to withdraw. The public function approve() sets the number of tokens (specified by the second parameter) that the caller’s address (i.e., the message sender) permits the first address parameter to withdraw. Therefore, ERC20 mandates that approve() updates the appropriate field of the two-dimensional mapping with the second parameter. To inject an error of this type into a function, we remove all assignments to the corresponding state variable within the function. The challenge is that ERCs do not prescribe how to name state variables, as these variables are only accessible within a contract. Consequently, given a rule of this type, different contracts are likely to use different names for the corresponding state variable. Fortunately, ERCs require a getter function to return its value for most state variables, and this getter function has the same name across all contracts for each ERC. This requirement enables us to locate the correct state variable automatically and perform the error injection.
@@
Violations of Function-Call Rules. Some ERC rules specify that a function must be called after a certain event. For instance, ERC721 requires yonERC721Received() to be called if tokens are sent to a contract. To inject an error for a call rule about function A() inside function B(), we simply remove all call sites of A() in B().
@@
Violations of Logging Rules. ERCs require specific events to be emitted within certain functions or after particular code actions for logging purposes. To inject such an error within a function, we remove all code statements that emit the corresponding event.
3.2 Dataset Summary

### v2

@@ Abstract
There is a huge demand to ensure the compliance of smart contracts listed on blockchain platforms to safety and economic standards described in natural languages. Today, manual efforts in the form of auditing are commonly used to achieve this goal. ML-based automated techniques have the promise to alleviate human efforts and the resulting monetary costs. However, unlike other domains where ML techniques have had huge successes, no systematic ML techniques have been proposed or applied to smart contract auditing. We present SC-Bench, the first dataset for automated smart-contract auditing research. SC-Bench consists of 5,377 real-world smart contracts running on Ethereum, a widely used blockchain platform, and 15,975 violations of standards on Ehereum called ERCs. Out of these violations, 139 are real violations programmers made. The remaining are errors systematically injected by us to reflect the violations of different ERC rules. We evaluate SC-Bench using GPT-4 by prompting it with both the contracts and ERC rules. In addition, we manually identify each violated rule and the corresponding code site (i.e., oracle) and prompt GPT-4 with the information asking for a True-or-False question. Our results show that without the oracle, GPT-4 can only detect 0.9% violations, and with the oracle, it detects 22.9% violations. These results show the potential room for improvement in ML-based techniques for smart-contract auditing.
Index Terms:
Smart Contract Auditing, Dataset
I Introduction
@@
Ethereum [35, 13] is a decentralized, open-source blockchain platform that has become the de facto for running decentralized applications like smart contracts [14, 17]. To standardize smart contract implementation, Ethereum Request for Comments (ERCs) have been developed. Each ERC provides a set of formal standards and is typically written in natural languages [31]. For example, ERC20 [33] defines the rules for fungible tokes — digital assets that are interchangeable. ERCs are essential in the Ethereum ecosystem, offering a common framework that developers must follow when implementing smart contracts. Violations of ERC rules can result in interoperability issues, contract failures, and financial loss. Moreover, tokens that don’t comply with ERCs may be delisted from exchanges, as many exchanges require ERC compliance for listing [15].
Despite the importance of adhering to ERC standards, developers often find it challenging to comply fully due to the complexity of the rules and their contract code. ERC standards consist of numerous rules, and this number continues to grow as new standards are introduced. For the three ERCs discussed in this work, there are 132 rules. These rules are presented in various formats, with some outlined as code comments and others explained in natural language paragraphs. Meanwhile, smart contract code is typically intricate, often spanning thousands of lines across multiple files. Some details may be hidden within complex caller-callee relationships, while others might involve objects and functions coded by different developers. This combination of ERC complexity and the intricacies of smart contracts makes it exceedingly difficult for programmers to ensure full compliance with ERC rules. Consequently, ERC violations are common in real-world smart contracts [8].
@@
To detect ERC violations, today’s common practice heavily relies on human efforts. Automated, program-analysis-based checkers for ERC rules do exist [9, 24], but they fail to detect complex violations because of many ERC rules’ non-structured, natural-language-based definitions. As a result, smart contracts commonly undergo human auditing, provided by specialized services with security experts [6, 28, 25, 3, 20, 1, 8]. Auditing services are not only costly but also slow. For example, we examined the history of 30 smart contracts that were submitted for manual auditing on a platform called Ethereum Commonwealth Security Department [8]. We found that each contract only has an average of 260 lines of code but is audited for ten days with an estimated cost of \$500. Clearly, manual auditing is not a scalable approach.
@@
A promising, scalable approach for automated smart-contract auditing is to leverage large language models (LLMs), both because of LLMs’ success in domains like program generation [30] and bug fixing [21, 38] and because of a fair amount of ERC rules’ natural-language descriptions. To develop LLM-based techniques (e.g., fine-tuning, few-shot learning) for smart contract auditing, an important step is to construct quality datasets. Unfortunately, no smart contract datasets exist for ERC rule checking and fixing. Traditional program bug datasets [29, 19] cannot be used for smart-contract auditing, as unlike ERC rules, compiler errors and program runtime errors have well-defined, structured definitions.
@@
To drive research and practices in smart-contract auditing and to assist real users with their auditing tasks, we release SC-Bench, the first dataset of real-world smart contracts and their ERC-rule violations. SC-Bench consists of 15,975 ERC violations and 5,377 real-world smart contracts, collected from etherscan.io [18], polygonscan.com [26], and Ethereum Commonwealth Security Department [8]. 139 violations from 30 contracts are real-world ERC violations we collect and inspect. As real violations are rarely published, we build program analysis techniques and inject 15,836 violations into 5,347 contracts according to 88 ERC rules.
@@
We evaluate SC-Bench using GPT-4 with two different approaches. First, we prompt GPT-4 with the contract to be inspected along with the full rule set of the corresponding ERC. Our results show that GPT-4 detects only 29% of real violations and 0.6% of injected violations. To assess how GPT-4 might perform with additional oracle information, we manually identify the violated ERC rules and the specific code sites causing the violations. We then prompt GPT-4 with this precise information and ask it a True-or-False question. In this case, GPT-4 successfully detects 42.8% of real violations and 22.8% of injected violations. The improved detection rates highlight the potential room for machine learning-based techniques in automating smart contract audits.
@@
Overall, SC-Bench goes beyond being a key contribution to Ethereum smart contract auditing. Other software auditing and verification tasks, such as ensuring the compliance of API usage rules, can potentially use SC-Bench’s contract dataset, its violation dataset, or its methodologies for evaluation.
We have released our dataset, results, and source code of our injecting scripts, all of which can be found at https://github.com/charlesxsh/scbench.
II Background
@@
II-A Ethereum and Smart Contracts
@@
II-B Ethereum Request for Comment (ERC)
@@ II-C ERC Rule Violations
TABLE I: How ERC rules are distributed across different error-injection methods. (“Uncovered”: rules whose violations cannot be injected by our designed error-injection methods.)
@@
II-D Today’s Auditing Practices
The common practice for detecting ERC rule violations today relies on manual auditing, often provided by paid services [6, 28, 25, 3, 20, 1, 8]. One such service is the Ethereum Commonwealth Security Department [8], where users submit smart contracts for auditing by filing a GitHub issue. The service then manually audits the submitted contracts and provides feedback through the issue.
To reduce the manual workload and associated costs, some automated tools have been developed using static program analysis. For example, Slither offers specific checkers (i.e., slither-check-erc [9]) that verify whether a given contract complies with the corresponding ERC standards for 11 ERCs. However, these tools have limited functionality. They primarily focus on ensuring the presence of required functions and events, confirming that these elements are correctly declared, and verifying that functions trigger the necessary events. Unfortunately, they are unable to check more advanced conditions, such as verifying whether the message caller has enough privilege to transfer tokens. To our knowledge, no machine learning-based automated auditing techniques have been proposed yet. We hope that the release of SC-Bench can foster a new line of research in this regard.
TABLE II: How many errors are injected by different error-injection methods.
# of Violations # of Cont.
@@ II-D Today’s Auditing Practices
III SC-Bench
III-A Construction
We collect real-world smart contracts from Ethereum Commonwealth Security Department [8], etherscan.io [18], and polygonscan.com [26]. As the purpose of SC-Bench is for evaluating automated ERC-rule checking, we include ERC rule violations of these collected contracts in SC-Bench using two approaches. First, we manually inspect a set of real smart contracts and identify all their ERC violations. This process is time-consuming, resulting in a limited number of violations. To address this, we perform automated error injection into a large number of real-world smart contracts, significantly increasing the number of violations. These two types of violations serve to validate each other and help ensure that the evaluated techniques demonstrate consistent performance.
Manual Inspection. We manually analyze 30 ERC20 contracts obtained from the Ethereum Commonwealth Security Department [8], selecting the most recent 30 audit requests that meet the following criteria: 1) they contain Solidity source code, 2) they have been approved by Solidity developers (as indicated by the “approved” tag), 3) they exhibit ERC rule violations, and 4) all contracts and the related Solidity code are within the same contract file. On average, each contract file contains 260.9 lines of Solidity code. Through detailed examination, we identify 139 ERC rule violations. Of these, 27 violations present a clear method for exploitation that can lead to financial losses, and we classify them as having a high-security impact. Another 48 violations result from incorrect implementation of required functionalities, but there is no apparent way to exploit these for financial gain. Therefore, we consider them to have a medium-security impact. The remaining 64 violations involve failures in generating necessary logs and are categorized as having a low-security impact.
1 contract ERC1155 {
2 function safeBatchTransferFrom(address from, address to, uint256[] memory ids, uint256[] memory amounts, bytes memory data) public {
3 require(from==_msgSender()||isApprovedForAll(from, _msgSender()), "not token owner or approved");
4 _safeBatchTransferFrom(from,to,ids,amounts,data);
5 }
6
7 function _safeBatchTransferFrom(address from, address to, uint256[] memory ids, uint256[] memory amounts, bytes memory data) internal {
8 require(ids.length==amounts.length, "ids and amounts length mismatch");
9- require(to!=address(0), "transfer to address 0");
10 address operator = _msgSender();
11 for (uint256 i = 0; i < ids.length; ++i) {
12 uint256 id = ids[I];
13 uint256 amount = amounts[I];
14 _balances[id][from] -= amount;
15 _balances[id][to] += amount;
16 }
17 emit TransferBatch(operator,from,to,ids,amounts);
18 }
19}
Figure 2: Violation injection of a condition-check rule. (Line 9 is deleted to perform the violation injection.)
Error Injection. To augment the limited manual-inspected real violations, we perform error injection to 5,347 real-world smart contract source code using the following methodology. We design six error injection methods corresponding to 85 ERC rules, resulting in a total of 15,836 ERC violations. Table I shows the number of rules each method covers.
We focus on contracts in three ERC standards, ERC20 [33], ERC721 [12], and ERC1155 [27], for two reasons. First, these ERCs are significant and have numerous crucial financial applications. For instance, there are over 450,000 ERC20 tokens on the Ethereum platform [5], many of which have market capitalizations exceeding \$1 billion (e.g., USDT [37], SHIB [36], Binance USD [34]). Second, these three standards are among the most mature ERCs and inspire subsequent ERCs, making their rules representative of rules in other ERCs. As shown in Table I, these three ERCs specify 132 rules. Our error injection methods can cover 85 of them. We do not include the rest either because they are not clearly specified (e.g., “throw if any other error” in ERC1155) or because they require more complex static analysis or injection methods.
@@ III-A Construction
We collect contracts in these three ERCs from etherscan.io [18] and polygonscan.com [26]. These two platforms are the most popular analytics platforms for Ethereum and its sidechain Polygon [22], respectively. As shown in Table II, we collected 5,211 contracts that are supposed to target ERC20, 110 for ERC721, and 26 for ERC1155111We mainly rely on a contract’s name or the name of a base contract inherited by the contract to determine whether the contract implements a particular ERC. . On average, each of these contracts contains 477.5 lines of code. We collect more contracts targeting ERC20 than ERC721 and ERC1155 contracts due to the significantly higher prevalence of ERC20 in practice.
@@
We randomly inject one to three errors in each contract based on the fit of rules to the contract to construct ERC violations. To inject an error, we first convert each input smart contract source-code file into its abstract syntax tree (AST). We then randomly select a rule and apply the corresponding error-injection method to the AST. After modifying the AST, we convert it back into source code. We perform error injections on ASTs because it is easier to conduct static program analysis and modify code on ASTs than to work directly with the source code. Table II shows the number of errors injected by each method. Finally, we verify that the modified source code compiles without errors using the Solidity compiler. Below, we discuss each error injection method in detail.
1contract ERC20 {
2 mapping(address => uint256) _balances;
3
4 function balanceOf(address account) public view returns (uint256) {
5- return _balances[account];
6+ return _balances[account]+827;
7 }
8}
Figure 3: Violation injection of a return rule. (Line 5 is replaced with line 6 to perform the violation injection.)
@@
Violations of Condition-Check Rules. ERCs mandate certain public functions to perform condition checks on their input parameters or callers’ addresses (i.e., msg.sender). Sometimes, these checks validate the input parameters. For example, ERC721 disallows the input parameter of function ownerOf() to be zero. More crucially, these checks ensure the caller has the permission to perform an action. For instance, ERC20’s transfer(address _to, uint256 _value) function sends tokens in the amount specified by its second parameter from the caller’s account to the receiver, whose address is the first parameter. ERC20 requires verifying that the caller has enough tokens (an amount greater than or equal to the second parameter) and throwing an exception if this condition is not met.
Many of these rules are enforced using require statements. To inject an error of this type into a function, we remove all require statements that check the parameter required by the rule as part of their conditions. Figure 2 illustrates an example of how a rule violation is injected in this way. ERC1155 requires that function safeBatchTransferFrom() “MUST revert if ‘_to’ is the zero address,” where _to is the second parameter of the function. Violating this rule would result in a permanent loss of the transferred tokens. The violation injection method begins with the safeBatchTransferFrom() function in line 2, as the rule applies to this function. It ignores the require statement in line 3, since its conditions do not involve the _to parameter. The method then proceeds to analyze the callee function _safeBatchTransferFrom() defined in lines 7–18. Similar to line 3, the method retains the require statement in line 7. However, it removes the require statement in line 9, as this one checks the _to parameter. None of the remaining lines contain require statements, so they are left unchanged. As shown in Figure 2, the modified contract can still be compiled by the Solidity compiler, but it allows safeBatchTransferFrom() to transfer tokens to the zero address, thus violating the rule.
@@
Violations of Return-Value Rules. ERCs explicitly define how return values should be computed for certain function call APIs. For example, ERC20 requires that allowance(address _owner, address _spender) returns the token amount that _owner allows _spender to withdraw. As return-value rules apply to only four data types, integer, Boolean, address, and string, we design four distinct error-injection methods for each of them. We add a random integer to an integer return value, flip a Boolean return value, replace an address return with a random address, and change a string return value to an empty string if it is not already empty or replace an empty return string with a random string.
Figure 3 illustrates an example of injecting a violation of a rule that specifies how to generate a integer return value. According to ERC20, the function balanceOf(address _owner) must return “the account balance of the another account with address _owner.” The contract in Figure 3 adheres to this rule, as shown in line 5. To inject a violation, we modify the return value by adding a random number to it, as demonstrated in line 6.
@@
Violations of Value-Update Rules. Fields or state variables in a contract represent its state. Some ERC rules specify how a public function should update a state variable. For example. contracts following ERC20 rules use a two-dimensional mapping to record the number of tokens the first key allows the second key to withdraw. The public function approve(address _spender, uint256 _value) sets the number of tokens (specified by the second parameter _value) that the caller’s address (i.e., the message sender) permits the first address parameter _spender to withdraw. Therefore, ERC20 mandates that approve(address _spender, uint256 _value) updates the appropriate field of the two-dimensional mapping with _value. To inject an error of this type into a function, we remove all assignments to the corresponding state variable within the function. The challenge is that ERCs do not prescribe how to name state variables, as these variables are only accessible within a contract. Consequently, given a rule of this type, different contracts are likely to use different names for the corresponding state variable. Fortunately, ERCs require a getter function to return its value for most state variables, and this getter function has the same name across all contracts for each ERC. This requirement enables us to locate the correct state variable automatically and perform the error injection.
@@
Violations of Function-Call Rules. Some ERC rules specify that a function must be called after a certain event. For instance, ERC721 requires onERC721Received() to be called if tokens are sent to a contract. To inject an error for a call rule about function A() inside function B(), we simply remove all call sites of A() in B().
@@
Violations of Logging Rules. ERCs require specific events to be emitted within certain functions or after particular code actions for logging purposes. To inject such an error that violates one of those rules within a function, we remove all code statements within the function that emit the corresponding event.
III-B Dataset Summary